OpenStudy (anonymous):

If f(x) = sin x and g(x) = cos x, determine the following. f (4n + 1)(x)= g(4n + 1)(x)= f (4n + 2)(x)= g(4n + 2)(x)= f (4n + 3)(x)= g(4n + 3)(x)= f (4n + 4)(x)= g(4n + 4)(x)= Now, use your answers to calculate F^(157)(x) when F(x) = 4 sin x − 3 cos x. F(157)(x)=

myininaya (myininaya):

what does f(4n+1)(x) mean?

myininaya (myininaya):

is this meant to be \[f^{(4n+1)}(x) \text{ where this means the } (4n+1)th \text { derivative of} f\]

geerky42 (geerky42):

\[f(x)=\sin(x)\\f'(x)=\cos(x)\\f''(x)=-\sin(x)\\f'''(x)=-\cos(x)\\f''''(x)=\sin(x)=f(x)\] Notice how it loops, right? So here's the answer to first one: \(f^{(4n+1)}(x) = f'(x) = \cos(x)\) Is it clear? Can you handle the rest?